Outage: General DDoS on edpnet services

Dear customers,

We are currently experiencing an distributed denial of service attack (DDoS) on all of our services. This may cause you to experience problems with all of our services.  Our Network Operations team is investigating this outage.  Further updates will be released as soon as possible.

Please find all details below:

Start*: 15-09-2021 16:15
End*: 19-09-2021 22:24
Duration: //
Impacted services: All services are impacted
Affected Areas**: All areas

EDIT 16:44*: As soon as we identified the issue, we have started routing our data through a scrubbing service (which tries to identify harmful sources & data and removes it).

EDIT 17:19*: The situation seems to be improving as we identify more and more attack vectors and sources. It seems like for most impacted customers, the situation has improved markedly since approximately ~17h.

EDIT 16/09/2021 10:12*: During the night we had two more attacks. We are working with the authorities, who have confirmed they are looking into it and are doing everything in their power to find the responsible individuals. We were contacted by an individual who verified he was behind the attacks, asking for a ransom.

EDIT 13:18*: We are being DDoS’ed again. We are doing what we can to mitigate the issues.

EDIT 15:41*: The attacks are still on-going. To give you a perspective of what we are dealing with, this is what our filtering service has been able to filter out so far:

EDIT 16:39*: We are monitoring the situation closely, and are discussing the situation both with our anti-DDoS provider and Belgian official instances. We hope to improve the situation as soon as possible and will keep you in the loop of our progress.

EDIT 17/9 15:47*: The attacks have started again. All of our uplinks are entirely congested. We are trying to mitigate the DDoS once more.

EDIT 18/09 12:41*: Yesterday evening and past night we were hit with multiple attacks.  Currently no attacks since 03:00. We still expect more attacks to come.  We are monitoring 24/7.  And we will communicate accordingly.  We try to answer questions or comments on social media as good as we can.

EDIT 20/09 9:48: There were some attacks on Sunday afternoon. No new attacks since then. We keep monitoring and do our best to minimize impact in case of new spikes.

We apologize for the inconveniences.

Best regards
The edpnet team

*All times are listed in CET, Central European Time
** Check the different areas & zones on our support site


Outage: Datacenter Sint-Niklaas. SOLVED

Dear customers,

We are currently experiencing an outage with our datacenter in Sint-Niklaas, one of our core routers there has crashed.  Our Network Operations team is investigating this outage.  Further updates will be released as soon as possible.

EDIT 10:19*: The issues were resolved at 9h30. We are investigating further.

Please find all details below:

Start*: 13-08-2021 8:15*
End*: 13-08-2021 9:30*
Duration: Approx. 75 minutes
Impacted services: Mail/webmail services, my edpnet, etc.
Affected Areas**: ALL

 

We apologize for the inconveniences.

Best regards
The edpnet team

*All times are listed in CET, Central European Time
** Check the different areas & zones on our support site


Outage: Internet and fixed telephony – Walloon Brabant, Hainaut : SOLVED

Dear customers,

We are currently experiencing an outage in A71 area (Walloon Brabant, Hainaut: 010, 060, 064, 065, 067, 068, 069, 071). Our Network Operations team is investigating the outage.

Please find all details below:

Start*: 14-04-2021 11:58
End*: 14-04-2021 13:37 / 15h12
Duration: 1 hour 39 minutes
Impacted services: all internet services
Affected Areas**: A71: Walloon Brabant, Hainaut: 010, 060, 064, 065, 067, 068, 069, 071

 

We apologize for the inconveniences.

Best regards
The edpnet team

UPD 12h40: the problem has been located. Our engineers are working to get the link up as soon as possible.

UPD 13h13: we suspect broken card on top of the fiber cut. Our technician is on his way to replace the card. Estimated time of resolution : 15h00

UPD 13h40: the problem  on the main path (fiber cut) is fixed. The links are coming up.

UPD: 15h20: The faulty card has been replaced. The problem is fully resolved.


Outage: edpnet office numbers unreachable : SOLVED

Dear customers,

We have experienced an outage with our reachability by phone.  We were unable to receive or make any calls with our office numbers. Our Network Operations team has investigated and resolved the outage.

Please find all details below:

Start*: 22-03-2021 08:08
End*: 22-03-2021 09:40
Duration: 1 hour 32 minutes
Impacted services: edpnet support reachability
Affected Areas**: edpnet office

 

We apologize for the inconveniences.

Best regards
The edpnet team

*All times are listed in CET, Central European Time
** Check the different areas & zones on our support site


Outage: problem with mobile telephony : SOLVED

Dear customers,

We have experienced an outage with mobile telephony.

Please find all details below:

Start*: 18-03-2021 14:20
End*: 18-03-2021 14:35
Duration: 15 minutes
Impacted services: mobile telephony
Affected Areas**: all areas

 

We apologize for the inconveniences.

Best regards
The edpnet team

*All times are listed in CET, Central European Time
** Check the different areas & zones on our support site


Outage: problem with mobile data, 4G/LTE backup : SOLVED

Dear customers,

We have experienced an outage with mobile data, 4G & LTE backup services.

Please find all details below:

Start*: 19-01-2021 11:30
End*: 19-01-2021 14:10
Duration: 2 hours 40 minutes
Impacted services: mobile data, 4G/LTE backup
Affected Areas**: all areas

 

We apologize for the inconveniences.

Best regards
The edpnet team

*All times are listed in CET, Central European Time
** Check the different areas & zones on our support site


Outage: High latency and speed issues – region: Liège, Namur, province of Luxembourg : SOLVED

Dear customers,

Our fiber connection between Brussels and Liege has been cut. All traffic has been rerouted.

Customers may notice higher latency and speed issues.

Please find all details below:

Start*: 2020-11-13 08:58
End*: 2020-11-13 10:55
Duration: 1h57min
Impacted services: high latency, speed issues
Affected Areas**: 41GRE: Liège, Namur, province of Luxembourg

UPDATE 10h44: technicians are currently fixing the faulty fiber joint

UPDATE 11h00: the link is restored. Traffic flow is coming back to normal.

We apologize for the inconveniences.

Best regards
The edpnet team

*All times are listed in CET, Central European Time
** Check the different areas & zones on our support site


Outage : Problem with internet services – SOLVED

Dear customers,

We are currently experiencing an outage with internet services (traffic drops). Our Network Operations team is investigating the outage.

Further details on the affected areas and an estimated duration will be released as soon as possible.

EDIT 19:40*: Our DNS-servers were the target of a concerted DDOS-attack for approximately one hour (18:00 – 19:00), involving large amounts of traffic and major network saturation. At this time the attack has cleared up and we are investigating further.

EDIT 30/8 20:35*: It seems like we’re being attacked again since a few minutes. Our Network Operations team is investigating. More updates will follow soon.

EDIT 20:45*: The attack was mitigated after some actions from our Network Operations team. It was smaller than last time, but still very much able to knock out our DNS-servers. We’re working to restore all services.

EDIT: 21:15*: All services are restored.

EDIT: 31/8 5h05 We experienced other waves of DDoS attacks between 30/8 22h20 and 31/08 05h05. For the moment, the attacks are fully mitigated.

We apologize for the inconveniences.

Best regards
The edpnet team


Outage notification: Internet services down – regions: Ath & Châtelet : SOLVED

UPDATE 11h45: The issue is also solved for Ath

UPDATE 11h25: Outage in Châtelet has been solved.

Dear customers,

We are currently experiencing an outage with internet services in regions Ath and Châtelet. Our Network Operations team is investigating the outage.

Further details on the affected areas and an estimated duration will be released as soon as possible.

We apologize for the inconveniences.

Best regards
The edpnet team


Outage: Internet services down : SOLVED

UPDATE 14h15: all links have been restored. The problem is completely resolved. Further details about this major incident will be shared later this week.

UPDATE 13h05: main path is restored , the peerings are being re-enabled , the traffic flow should get back to normal very soon.

Dear customers,

We are currently experiencing a fiber cut within our network. Internet and telephony services are unavailable for our customers in the Netherlands. Our Network Operations team is investigating the outage.

Further details  and an estimated duration will be released as soon as possible.

We apologize for the inconveniences.

Best regards
The edpnet team